Find the magnitude of the gravitational forceon Pluto, with a mass of 1.32 × 1022kg and aradius of 1.15 × 106 m.Answer in units of N
Answer #1
m1 = 63.5 kg
m2 = 5.98 *10^24 kg
r = 6.37 *10^6 m
G = 6.673 × 10-11 Nm2/kg2
Since F = Gm1m2/r^2
F =(6.673 × 10-11 Nm2/kg2) * 63.5 kg * 5.98 *10^24 kg / (6.37 *10^6 m)^2
F = 624.4 N

please post each question separately.
answered by: ballamoney1
Answer #2
Gravitational force experienced by him on Pluto =6.67*10^-11 * 6.63*10 * 1.32*10^22 / (1.15*10^6)^2 = 44.2 N
answered by: Kaitlynn
Answer #3

For any mass in universe there exists a gravtational pull on any other mass,
this field varies over the distance from the centre of that mass can be calculated using
g=GM/r^2; G=6.673*10^-11Nm2/kg2.
mass of pluto is 1.32*10^22kg,radius=1.15*106m;
g=0.66m/s2;
is the acceleration due to gravity on the surface of pluto.
The Gravitationalforce is F=GMm/r^2;
where M is the mass of planet(here pluto),m is the mass of the body on which we are aiming to calculate,here we need the force of gravtation on plutohence by taking distance between the bodies as the radius we'll get
F=0.66*(m)N.m is the mass of body.

therefore F=41.91N.

answered by: justme